# NOCTIS: Novel Object Cyclic Threshold based Instance Segmentation

### v2

@@ Abstract
Instance segmentation of novel objects instances in RGB images, given some example images for each object, is a well known problem in computer vision. Designing a model general enough to be employed for all kinds of novel objects without (re-) training has proven to be a difficult task. To handle this, we present a new training-free framework, called: Novel Object Cyclic Threshold based Instance Segmentation (NOCTIS). NOCTIS integrates two pre-trained models: Grounded-SAM 2 for object proposals with precise bounding boxes and corresponding segmentation masks; and DINOv2 for robust class and patch embeddings, due to its zero-shot capabilities. Internally, the proposal-object matching is realized by determining an object matching score based on the similarity of the class embeddings and the average maximum similarity of the patch embeddings with a new cyclic thresholding (CT) mechanism that mitigates unstable matches caused by repetitive textures or visually similar patterns. Beyond CT, NOCTIS introduces: (i) an appearance score that is unaffected by object selection bias; (ii) the usage of the average confidence of the proposals’ bounding box and mask as a scoring component; and (iii) an RGB-only pipeline that performs even better than RGB-D ones. We empirically show that NOCTIS, without further training/fine tuning, attains state-of-the-art results regarding the mean AP score, w.r.t. the best RGB and RGB-D methods on the seven core datasets of the BOP 2023 challenge for the “Model-based 2D segmentation of unseen objects” task.  [^1] Code available at: https://github.com/code-iai/noctis
@@ 1 Introduction
Instance segmentation (and detection) is one of the key problems in robot perception and augmented reality applications, since it tries to identify and locate object instances in images or videos via bounding boxes and segmentation masks; e.g. a robot wants to identify a specific object instance on a conveyor belt. Classical supervised deep learning methodologies such as Ren et al. (2016); He et al. (2018); Lin et al. (2018); Labbé et al. (2020); Su et al. (2022) have achieved good performances, however, for each target object they require an expensive (re-) training or, at least, a fine-tuning step. Hence, most of the time, adding novel/unseen objects begs for additional training data, either synthetic or real, with (human-made) annotations. As a consequence, the usage of said supervised methods for industry, especially for short-time prototype development cycles, where the target object changes constantly, is unfeasible.
In this work, we introduce NOCTIS, a fully training-free RGB-only framework, for novel objects instance segmentation that achieves state-of-the-art (SOTA) performances on the BOP 2023 core datasets. NOCTIS builds upon the strengths of two recent pre-trained models, Grounded-SAM 2 (Ren et al., 2024b) (GSAM 2) for precise mask proposal generation and DINOv2 (Oquab et al., 2024) for robust visual descriptors, but departs from previous approaches like CNOS (Nguyen et al., 2023); NIDS-Net (Lu et al., 2025) and SAM-6D(Lin et al., 2024) in several critical ways.
@@
Second, we propose a cyclic thresholding (CT) mechanism, a novel patch-filtering strategy designed to handle repetitive textures and visually similar patterns that can lead to many‑to‑one matches. Unlike nearest neighbor matching (e.g. Simakov et al. (2008); Oron et al. (2018)), CT relaxes strict mutual matching requirements, tolerating some distance between a patch and its cyclic/round-trip patch, while filtering out unreasonable ones (see Section 3.3). This yields a more reliable proposal-template matching and improves the appearance-based scoring process.
Third, we are the first, to the best of our knowledge, to incorporate mask and bounding-box confidence values, readily available from modern mask proposal generators, into the final object matching score. While such confidence measures are commonly produced by detection and segmentation models (e.g. Grounding-DINO (Liu et al., 2024); SAM (Kirillov et al., 2023); FastSAM (Zhao et al., 2023); Grounded-SAM (Ren et al., 2024b) (GSAM)), they have not been exploited in this task before, and we demonstrate their positive impact through ablation studies.
Finally, we test our approach on the seven core datasets of the BOP 2023 challenge (Hodan et al., 2024) for the “2D instance segmentation of unseen objects” task; and we show that our method NOCTIS, without further training, performs better than other RGB and RGB-D methods in terms of the Average Precision (AP) metric; challenging the assumption that depth information is required for top-tier performances in this domain. Moreover, we surpass the best published method NIDS-Net (Lu et al., 2025) by a significant margin of absolute $3.4\%$ mean AP; while for the unpublished ones, we are on par with the (updated) best one and overcoming the second best one by $0.8\%$.
@@
  • We propose NOCTIS, an RGB-only zero-shot novel objects instance segmentation framework that uses foundation models and performs on par or better than the SOTA ones.
  • An unbiased appearance score that aggregates over all templates to remove selection bias.
@@ 2 Related work
The usage of pre-trained foundation models has become pervasive due to their strong performance across diverse downstream tasks. Notably, research efforts such as Visual Transformers (ViT) (Dosovitskiy et al., 2021); CLIP (Radford et al., 2021); DINOv2 and others (Caron et al., 2021; Cherti et al., 2023); have focused on large-scale image representation learning to improve generalization. These models encapsulate extensive visual knowledge, making them suitable as backbones for a variety of tasks, including image classification, video understanding, depth estimation, semantic segmentation, and novel instance retrieval. The main challenge, however, lies in harnessing their capabilities effectively for a specific target domain. We adopt DINOv2 as our feature extractor, leveraging its ability to produce high-quality and robust descriptors for previously unseen instances.
@@
Another area where foundation models currently excel is image segmentation/semantic mask generation, with Segment Anything (SAM), a ViT-based model, being the forerunner. Since SAM can be computationally demanding, several variations have been proposed for real-world scenarios to reduce costs by replacing components with smaller (less parameters/weights) ViT models (Zhang et al., 2023; Ke et al., 2023), or even CNN-based ones (Zhao et al., 2023; Zhou et al., 2024; Wang et al., 2024). Notably, its successor, SAM 2 (Ravi et al., 2024), while having some additional features like video tracking, achieves a higher quality in terms of Mean Intersection over Union (mIoU) than SAM and is also more efficient computation- and memory-wise. In recent times, it has become an established practice to combine the strengths of multiple models in a modular way to solve complex problems. Indeed, a standard practice to tackle segmentation problems consists of combining open-set object detectors (Li et al., 2022; Jiang et al., 2024; Ren et al., 2024a; Liu et al., 2024) with a SAM variant. We adopt GSAM 2 in this modular spirit but extend its utility beyond simple mask generation by incorporating its bounding box and mask confidence values directly into our scoring framework; an element absent in prior work.
@@
Traditionally, instance segmentation methods, like Mask R-CNN (He et al., 2018) or similar (Ren et al., 2016; Lin et al., 2018; Su et al., 2022), used to be fine-tuned on specific target objects (Sundermeyer et al., 2023). Even though these methods have been demonstrated to be robust in challenging scenarios with heavy occlusions and lighting conditions, they lack the flexibility to handle novel objects without retraining, known as a “closed-world” setting; a limitation that significantly hinders their applicability in real-world settings. To overcome this limitation, progress was made in the task of novel object instance segmentation, where ZeroPose (Chen et al., 2024) and CNOS (Nguyen et al., 2023) were among the first notable ones solving it. The core architecture of the latter, using template views as references and a SAM variant for proposal creation while classifying these via a similarity-based image matching technique, has laid the foundation for subsequent models such as SAM-6D and NIDS-Net. These works inspired us to adopt said approach as our starting point.
@@ 3 Method
Our approach, as shown in Figure 1, is carried out in three steps. Starting with the onboarding stage in Section 3.1, visual descriptors are extracted from the template images via DINOv2; followed by the proposal stage in Section 3.2, where all possible segmentation masks and their descriptors, from the query RGB image, are generated with GSAM 2 and DINOv2, respectively. Lastly, in Section 3.3, the matching stage, each proposed mask is given an object label and a confidence value, based on the determined object scores using the visual descriptors.
Figure 1: The three NOCTIS stages: onboarding stage, represents each object via descriptors from templates (Section 3.1); proposal stage (Section 3.2), where proposals (masks) and their descriptors from the query RGB image are generated; lastly, in the matching stage, object labels and confidences are assigned to each proposal based on their descriptors (Section 3.3).
@@ 3.1 Onboarding stage
The goal of the onboarding stage is to generate multiple visual descriptors to represent each of the $N^{O}$ different novel objects $\mathbb{O}$}. In the following, in all the descriptions and notations, we will consider just one object $\mathit{O}\in{\mathbb{O}}$; this is done to keep the notation simple. In detail, the object $\mathit{O}$ is represented by a set of $N^{T}$ template images $\mathbb{T}$} (${{\mathbb{R}}}^{3\times\text{W}^{\prime}\times\text{H}^{\prime}}$ images) and their corresponding ground truth segmentation masks showing the object from different predefined viewpoints. Given some fixed viewpoints, there are multiple possible sources for these templates and masks, e.g. pre-render them with renderers like Pyrender (Matthew Matl, 2021) or BlenderProc (Denninger et al., 2023); or even extract them out of some selected frames, e.g. annotated videos, where the object is not too occluded and has a viewpoint close to a predefined one.
@@ 3.2 Proposal stage
At this stage, all object proposals from the query image $\bm{I}$} are acquired. While previous works (Li et al., 2023; Shen et al., 2023; Chen et al., 2024; Nguyen et al., 2023; Lin et al., 2024) have employed various “pure” SAM-based proposal generators, we instead adopt GSAM 2 as a modified version of GSAM used in NIDS-Net. The original GSAM obtains the bounding boxes of all objects from Grounding-DINO (Liu et al., 2024), a pre-trained zero-shot detector, matching a given text prompt; then, it uses these as a prompt for SAM to create masks. GSAM 2 replaced its SAM component with the qualitative (w.r.t. mIoU) and performance-wise improved SAM 2.
@@ 3.3 Matching stage
with $\langle,\rangle$ denoting the inner product and $\|\cdot\|$ the Euclidean norm. If the vectors point in the same direction, they have a $\mathit{cossim}$ value of $1$, 1-1 for opposite directions and 0 for orthogonality. It was shown in CNOS (Nguyen et al., 2023, Section 4.3), that for this score, using the top-$5$ average as an aggregating function, is the most robust option out of: Mean; Max; Median and top-K Average.
@@
$\mathit{s}_{\mathit{p},{\bm{T}}}^{\mathit{appe}}=\frac{1}{N_{{\bm{I}}_{\mathit{p}}}^{\mathit{crop}}}\sum_{i=1}^{N_{{\bm{I}}_{\mathit{p}}}^{\mathit{crop}}}\max_{j=1,\dots,N_{{\bm{T}}}^{\mathit{crop}}}\big{(}\mathit{cossim}({{\bm{F}}}_{{{\bm{I}}}_{\mathit{p}},:,i}^{\mathit{patch}},{{\bm{F}}}_{{\bm{T}},:,j}^{\mathit{patch}})\big{)}\cdot\bm{1}_{\mathit{cdist}({{\bm{I}}}_{\mathit{p}},{\bm{T}},i)\leq\delta_{\mathit{CT}}}\text{,}$ (2)
where ${{\bm{F}}}_{{{\bm{I}}}_{\mathit{p}},:,i}^{\mathit{patch}}$ represents the $i$-th column of the ${{\bm{F}}}_{{{\bm{I}}}_{\mathit{p}}}^{\mathit{patch}}$ patch token matrix; ${{\bm{F}}}_{{\bm{T}},:,j}^{\mathit{patch}}$ the $j$-th column of ${{\bm{F}}}_{{\bm{T}}}^{\mathit{patch}}$; and $\bm{1}$ the indicator function, which is used to filter out certain patch pairs. As these sub-appearance scores show an inherent sensitivity to the visible parts of the object instances, thus to the viewpoint differences, they are aggregated into the final appearance score $\mathit{s}_{\mathit{p}}^{\mathit{appe}}$ using the Max function (across the templates), to mitigate this phenomenon.
@@
The idea behind the cyclic threshold (CT) filtering arises as DINOv2 descriptors can assign similar patch tokens to repetitive textures/similar looking parts (e.g. identical corners or surfaces), leading to many‑to‑one matches, which one would like to avoid via patch filtering of sorts. An often used technique to solve this issue is the nearest neighbor based image patch matching, e.g. for finding the most coherent pairs of patches Simakov et al. (2008) and points Oron et al. (2018); however, in practice, this incurs in a restrictive filtering, that is why we relaxed this bidirectional similarity aspect to account for a non strictly one-to-one mapping assumption between the template and the query regions, as different scene lighting and occlusions might dampen it. Therefore, our CT filtering allows a matching that is not just strictly mutual but permits a certain degree of tolerance, i.e. how many patches in the neighborhood of the considered ones, in terms of Euclidean distance, are still accepted.
In Figure 2, the general representation of the cyclic distance of patch $s$, through $t$ and $u$, is given; where $s$ and $u$ belong to image $\bm{A}$} and $t$ to $\bm{B}$}. Each example image is divided into an $8\times 8$ grid for the sake of simplicity, resulting in 64 patches. From patch $s$ in image $\bm{A}$}, we find $t$, the most similar patch of it in $\bm{B}$}, via the following function:
$\mathit{bestMatchIndex}({{\bm{F}}}_{{\bm{A}}}^{\mathit{patch}},{{\bm{F}}}_{{\bm{B}}}^{\mathit{patch}},i)=\underset{j=1,\dots,N_{{\bm{B}}}^{\mathit{crop}}}{\mathrm{argmax}}\mathit{cossim}({{\bm{F}}}_{{\bm{A}},:,i}^{\mathit{patch}},{{\bm{F}}}_{{\bm{B}},:,j}^{\mathit{patch}})\text{,}$ (3)
@@
Given the previous discussion regarding the mutual similarity principle, using a CT value of 0 is equivalent to mutual nearest neighbor based image patch matching. Therefore using the previously defined cyclic distance, our addition to the appearance score lies in the application of a patch filter, represented by the indicator function $\bm{1}$ in equation 2, to increase the score’s reliability/expressiveness; which allows for a relaxed mutual nearest neighbours matching. The function $\bm{1}_{\mathit{cdist}({{\bm{I}}}_{\mathit{p}},{\bm{T}},i)\leq\delta_{\mathit{CT}}}$ internally calculates the $\mathit{cdist}$ for patch $i$ of image crop ${{\bm{I}}}_{\mathit{p}}$ and template $\bm{T}$}, then checks if it is smaller than a predefined CT value. The default value for it is $\delta_{\mathit{CT}}=5$, see also the Appendix Section A.1, where the effects of using different threshold values are discussed.
Lastly, to sum up our contributions to the appearance score, one can see that it significantly differs from the approach used in SAM-6D (Lin et al., 2024, Section 3.1.2), as its authors computed only the sub-appearance score for the single template of the object having the highest semantic score, thus resulting in a highly biased score. Indeed, as it was shown in CNOS (Nguyen et al., 2023, Section 4.4), the cls token contains insufficient information about matching viewpoints, potentially leading to low appearance values. Additionally, we included our CT filtering technique to improve the appearance score accuracy even further.
@@
$\mathit{s}_{\mathit{p}}^{\mathit{obj}}=\frac{\mathit{s}_{\mathit{p}}^{\mathit{sem}}+w_{\mathit{appe}}*\mathit{s}_{\mathit{p}}^{\mathit{appe}}}{1+1}*\mathit{conf}_{\mathit{p}}\text{,}$ (4)
where an appearance weight of $w_{\mathit{appe}}=1$ computes the average. The object matching scores of all the $N_{{\bm{I}}}^{\mathit{prop}}$ proposals, over all possible $N^{O}$ objects, are stored in the $N_{{\bm{I}}}^{\mathit{prop}}\times N^{O}$ instance score matrix. Note that, as small CT values squash down the appearance scores, $w_{\mathit{appe}}=2$ is used.
@@
In the last step, we simply apply the Argmax function across the objects/rows of the instance score matrix. Each proposal gets assigned an object label and its object matching score as its corresponding confidence. Eventually, we obtain proposals consisting of: a bounding box of the object instance; its corresponding modal segmentation mask, which covers the visible instance part (Hodan et al., 2024); and an object label with a confidence score. To remove incorrectly labeled proposals and redundant ones, a confidence threshold filtering is applied with $\delta_{\mathit{conf}}=0.2$ followed by a Non-Maximum Suppression, respectively.

### v4

@@ Abstract
Instance segmentation of novel objects instances in RGB images, given some example images for each object, is a well known problem in computer vision. Designing a model general enough to be employed for all kinds of novel objects without (re-) training has proven to be a difficult task. To handle this, we present a new training-free framework, called: Novel Object Cyclic Threshold based Instance Segmentation (NOCTIS). NOCTIS integrates two pre-trained models: Grounded-SAM 2 for object proposals with precise bounding boxes and corresponding segmentation masks; and DINOv2 for robust class and patch embeddings, due to its zero-shot capabilities. Internally, the proposal-object matching is realized by determining an object matching score based on the similarity of the class embeddings and the average maximum similarity of the patch embeddings with a new cyclic thresholding (CT) mechanism that mitigates unstable matches caused by repetitive textures or visually similar patterns. Beyond CT, NOCTIS introduces: (i) an appearance score that is unaffected by object selection bias; (ii) the usage of the average confidence of the proposals’ bounding box and mask as a scoring component; and (iii) an RGB-only pipeline that performs even better than RGB-D ones. We empirically show that NOCTIS, without further training/fine tuning, outperforms the best RGB and RGB-D methods regarding the mean AP score on the seven core datasets of the BOP 2023 challenge for the “Model-based 2D segmentation of unseen objects” task.  [^1] Code available at: https://github.com/code-iai/noctis
@@ 1 Introduction
Instance segmentation (and detection) is one of the key problems in robot perception and augmented reality applications, since it tries to identify and locate object instances in images or videos via bounding boxes and segmentation masks; e.g. a robot wants to identify a specific object instance on a conveyor belt. Classical supervised deep learning methodologies such as ren2016fasterrcnnrealtimeobject, he2018maskrcnn, lin2018focallossdenseobject, labbe2020cosyposeconsistentmultiviewmultiobject, su2022zebraposecoarsefinesurface have achieved good performances, however, for each target object they require an expensive (re-) training or, at least, a fine-tuning step. Hence, most of the time, adding novel/unseen objects begs for additional training data, either synthetic or real, with (human-made) annotations. As a consequence, the usage of said supervised methods for industry, especially for short-time prototype development cycles, where the target object changes constantly, is unfeasible.
In this work, we introduce NOCTIS, a fully training-free RGB-only framework, for novel objects instance segmentation that achieves state-of-the-art (SOTA) performances on the BOP 2023 core datasets. NOCTIS builds upon the strengths of two recent pre-trained models, Grounded-SAM 2 (GSAM 2) [ren2024groundedsamassemblingopenworld] for precise mask proposal generation and DINOv2 [oquab2024dinov2learningrobustvisual] for robust visual descriptors, but departs from previous approaches like CNOS [nguyen2023cnosstrongbaselinecadbased]; NIDS-Net [lu2025adaptingpretrainedvisionmodels] and SAM-6D[Lin_2024_CVPR] in several critical ways.
@@
Second, we propose a cyclic thresholding (CT) mechanism, a novel patch-filtering strategy designed to handle repetitive textures and visually similar patterns that can lead to many‑to‑one matches. Unlike nearest neighbor matching ([4587842, 8006263]), CT relaxes strict mutual matching requirements, tolerating some distance between a patch and its cyclic/round-trip patch, while filtering out unreasonable ones (see Sec. 3.3). This yields a more reliable proposal-template matching and improves the appearance-based scoring process.
Third, we are the first, to the best of our knowledge, to incorporate mask and bounding-box confidence values, readily available from modern mask proposal generators, into the final object matching score. While such confidence measures are commonly produced by detection and segmentation models (Grounding-DINO [liu2024groundingdinomarryingdino]; SAM [kirillov2023segment]; FastSAM [zhao2023fastsegment]; Grounded-SAM (GSAM) [ren2024groundedsamassemblingopenworld]), they have not been exploited in this task before, and we demonstrate their positive impact through ablation studies.
Finally, we test our approach on the seven core datasets of the BOP 2023 challenge [hodan2024bopchallenge2023detection] for the “2D instance segmentation of unseen objects” task; and we show that our method NOCTIS, without further training, performs better than other RGB and RGB-D methods in terms of the Average Precision (AP) metric; challenging the assumption that depth information is required for top-tier performances in this domain. Moreover, we surpass the best published method NIDS-Net [lu2025adaptingpretrainedvisionmodels] by a significant margin of absolute $3.4\%$ mean AP; while for the unpublished ones, we overcome the best among them by $0.8\%$.
@@
  • We propose NOCTIS, an RGB-only zero-shot novel objects instance segmentation framework that uses foundation models and performs better than the SOTA ones.
  • An appearance score that aggregates over all templates to remove selection bias.
@@ 2 Related work
The usage of pre-trained foundation models has become pervasive due to their strong performance across diverse downstream tasks. Notably, research efforts such as Visual Transformers (ViT) [dosovitskiy2021imageworth16x16words]; CLIP [radford2021learningtransferablevisualmodels]; DINOv2 and others [caron2021emergingpropertiesselfsupervisedvision, Cherti_2023]; have focused on large-scale image representation learning to improve generalization. These models encapsulate extensive visual knowledge, making them suitable as backbones for a variety of tasks, including image classification, video understanding, depth estimation, semantic segmentation, and novel instance retrieval. The main challenge, however, lies in harnessing their capabilities effectively for a specific target domain. We adopt DINOv2 as our feature extractor, leveraging its ability to produce high-quality and robust descriptors for previously unseen instances.
@@
Another area where foundation models currently excel is image segmentation/semantic mask generation, with Segment Anything (SAM), a ViT-based model, being the forerunner. Since SAM can be computationally demanding, several variations have been proposed for real-world scenarios to reduce costs by replacing components with smaller (less parameters/weights) ViT models [zhang2023fastersegmentanythinglightweight, ke2023segmenthighquality], or even CNN-based ones [zhao2023fastsegment, zhou2024edgesampromptintheloopdistillationondevice, wang2024repvitsamrealtimesegmenting]. Notably, its successor, SAM 2 [ravi2024sam2segmentimages], while having some additional features like video tracking, achieves a higher quality in terms of Mean Intersection over Union (mIoU) than SAM and is also more efficient computation- and memory-wise. In recent times, it has become an established practice to combine the strengths of multiple models in a modular way to solve complex problems. Indeed, a standard practice to tackle segmentation problems consists of combining open-set object detectors [li2022groundedlanguageimagepretraining, jiang2024trex2genericobjectdetection, ren2024groundingdino15advance, liu2024groundingdinomarryingdino] with a SAM variant. We adopt GSAM 2 in this modular spirit but extend its utility beyond simple mask generation by incorporating its bounding box and mask confidence values directly into our scoring framework; an element absent in prior work.
@@
Traditionally, instance segmentation methods, like Mask R-CNN [he2018maskrcnn] or similar [ren2016fasterrcnnrealtimeobject, lin2018focallossdenseobject, su2022zebraposecoarsefinesurface], used to be fine-tuned on specific target objects [sundermeyer2023bopchallenge2022detection]. Even though these methods have been demonstrated to be robust in challenging scenarios with heavy occlusions and lighting conditions, they lack the flexibility to handle novel objects without retraining, known as a “closed-world” setting; a limitation that significantly hinders their applicability in real-world settings. To overcome this limitation, progress was made in the task of novel object instance segmentation, where ZeroPose [chen2024zeroposecadpromptedzeroshotobject] and CNOS [nguyen2023cnosstrongbaselinecadbased] were among the first notable ones solving it. The core architecture of the latter, using template views as references and a SAM variant for proposal creation while classifying these via a similarity-based image matching technique, has laid the foundation for subsequent models such as SAM-6D and NIDS-Net. These works inspired us to adopt said approach as our starting point.
@@ 3 Method
Our approach, as shown in Fig. 1, is carried out in three steps. Starting with the onboarding stage in Sec. 3.1, visual descriptors are extracted from the template images via DINOv2; followed by the proposal stage in Sec. 3.2, where all possible segmentation masks and their descriptors, from the query RGB image, are generated with GSAM 2 and DINOv2, respectively. Lastly, in Sec. 3.3, the matching stage, each proposed mask is given an object label and a confidence value, based on the determined object scores using the visual descriptors.
Figure 1: The three NOCTIS stages: onboarding stage, represents each object via descriptors from templates ( Sec. 3.1); proposal stage ( Sec. 3.2), where proposals (masks) and their descriptors from the query RGB image are generated; lastly, in the matching stage, object labels and confidences are assigned to each proposal based on their descriptors ( Sec. 3.3).
@@ 3.1 Onboarding stage
The goal of the onboarding stage is to generate multiple visual descriptors to represent each of the $N^{O}$ different novel objects $\mathbb{O}$}. In the following, in all the descriptions and notations, we will consider just one object $\mathit{O}\in{\mathbb{O}}$; this is done to keep the notation simple. In detail, the object $\mathit{O}$ is represented by a set of $N^{T}$ template images $\mathbb{T}$} (${{\mathbb{R}}}^{3\times\text{W}^{\prime}\times\text{H}^{\prime}}$ images) and their corresponding ground truth segmentation masks showing the object from different predefined viewpoints. Given some fixed viewpoints, there are multiple possible sources for these templates and masks, e.g. pre-render them with renderers like Pyrender [pyrender] or BlenderProc [Denninger2023]; or even extract them out of some selected frames, e.g. annotated videos, where the object is not too occluded and has a viewpoint close to a predefined one.
@@ 3.2 Proposal stage
At this stage, all object proposals from the query image $\bm{I}$} are acquired. While previous works [li2023voxdetvoxellearningnovel, shen2023highresolutiondatasetinstancedetection, chen2024zeroposecadpromptedzeroshotobject, nguyen2023cnosstrongbaselinecadbased, Lin_2024_CVPR] have employed various “pure” SAM-based proposal generators, we instead adopt GSAM 2 as a modified version of GSAM used in NIDS-Net. The original GSAM obtains the bounding boxes of all objects from Grounding-DINO [liu2024groundingdinomarryingdino], a pre-trained zero-shot detector, matching a given text prompt; then, it uses these as a prompt for SAM to create masks. GSAM 2 replaced its SAM component with the qualitative (w.r.t. mIoU) and performance-wise improved SAM 2.
@@ 3.3 Matching stage
with $\langle,\rangle$ denoting the inner product and $\|\cdot\|$ the Euclidean norm. If the vectors point in the same direction, they have a $\mathit{cossim}$ value of $1$, 1-1 for opposite directions and 0 for orthogonality. It was shown in CNOS [nguyen2023cnosstrongbaselinecadbased, Section 4.3], that for this score, using the top-$5$ average as an aggregating function, is the most robust option out of: Mean; Max; Median and top-K Average.
@@
$\mathit{s}_{\mathit{p},{\bm{T}}}^{\mathit{appe}}=\Big(\sum_{i=1}^{N_{{\bm{I}}_{\mathit{p}}}^{\mathit{crop}}}$ $\max_{j=1,\dots,N_{{\bm{T}}}^{\mathit{crop}}}\big(\mathit{cossim}({{\bm{F}}}_{{{\bm{I}}}_{\mathit{p}},:,i}^{\mathit{patch}},{{\bm{F}}}_{{\bm{T}},:,j}^{\mathit{patch}})\big)$ (2)
$\cdot\bm{1}_{\mathit{cdist}({{\bm{I}}}_{\mathit{p}},{\bm{T}},i)\leq\delta_{\mathit{CT}}}\Big)\cdot\frac{1}{N_{{\bm{I}}_{\mathit{p}}}^{\mathit{crop}}}\text{,}$
where ${{\bm{F}}}_{{{\bm{I}}}_{\mathit{p}},:,i}^{\mathit{patch}}$ represents the $i$-th column of the patch token matrix ${{\bm{F}}}_{{{\bm{I}}}_{\mathit{p}}}^{\mathit{patch}}$; ${{\bm{F}}}_{{\bm{T}},:,j}^{\mathit{patch}}$ the $j$-th column of ${{\bm{F}}}_{{\bm{T}}}^{\mathit{patch}}$; and $\bm{1}$ the indicator function, which is used to filter out certain patch pairs. As these sub-appearance scores show an inherent sensitivity to the visible parts of the object instances, thus to the viewpoint differences, they are aggregated into the final appearance score $\mathit{s}_{\mathit{p}}^{\mathit{appe}}$ using the Max function (across the templates), to mitigate this phenomenon.
@@
The idea behind the cyclic threshold (CT) filtering arises as DINOv2 descriptors can assign similar patch tokens to repetitive textures/similar looking parts (e.g. identical corners or surfaces), leading to many‑to‑one matches, which one would like to avoid via patch filtering of sorts. An often used technique to solve this issue is the nearest neighbor based image patch matching, e.g. for finding the most coherent pairs of patches [4587842] and points [8006263]; however, in practice, this incurs in a restrictive filtering, that is why we relaxed this bidirectional similarity aspect to account for a non strictly one-to-one mapping assumption between the template and the query regions, as different scene lighting and occlusions might dampen it. Therefore, our CT filtering allows a matching that is not just strictly mutual but permits a certain degree of tolerance, i.e. how many patches in the neighborhood of the considered ones, in terms of Euclidean distance, are still accepted.
In Fig. 2, the general representation of the cyclic distance of patch $s$, through $t$ and $u$, is given; where $s$ and $u$ belong to image $\bm{A}$} and $t$ to $\bm{B}$}. Each example image is divided into an $8\times 8$ grid for the sake of simplicity, resulting in $64$ patches. From patch $s$ in image $\bm{A}$}, we find $t$, the most similar patch of it in $\bm{B}$}, via the following function:
$\mathit{bestMatchIndex}$ $({{\bm{F}}}_{{\bm{A}}}^{\mathit{patch}},{{\bm{F}}}_{{\bm{B}}}^{\mathit{patch}},i)=$ (3)
$=\underset{j=1,\dots,N_{{\bm{B}}}^{\mathit{crop}}}{\mathrm{argmax}}$ $\mathit{cossim}({{\bm{F}}}_{{\bm{A}},:,i}^{\mathit{patch}},{{\bm{F}}}_{{\bm{B}},:,j}^{\mathit{patch}})\text{,}$
@@
Given the previous discussion regarding the mutual similarity principle, using a CT value of 0 is equivalent to mutual nearest neighbor based image patch matching. Therefore using the previously defined cyclic distance, our addition to the appearance score lies in the application of a patch filter, represented by the indicator function $\bm{1}$ in Eq. 2, to increase the score’s reliability/expressiveness; which allows for a relaxed mutual nearest neighbours matching. The function $\bm{1}_{\mathit{cdist}({{\bm{I}}}_{\mathit{p}},{\bm{T}},i)\leq\delta_{\mathit{CT}}}$ internally calculates the $\mathit{cdist}$ for patch $i$ of image crop ${{\bm{I}}}_{\mathit{p}}$ and template $\bm{T}$}, then checks if it is smaller than a predefined CT value. The default value for it is $\delta_{\mathit{CT}}=5$, see also the Appendix Sec. 6.1, where the effects of using different threshold values are discussed.
Lastly, to sum up our contributions to the appearance score, one can see that it significantly differs from the approach used in SAM-6D [Lin_2024_CVPR, Section 3.1.2], as its authors computed only the sub-appearance score for the single template of the object having the highest semantic score, thus resulting in a highly biased score; while ours does not suffer from this kind of “bias”. Indeed, as it was shown in CNOS [nguyen2023cnosstrongbaselinecadbased, Section 4.4], the cls token contains insufficient information about matching viewpoints, potentially leading to low appearance values. Additionally, we included our CT filtering technique to improve the appearance score accuracy even further.
@@
$\mathit{s}_{\mathit{p}}^{\mathit{obj}}=\frac{\mathit{s}_{\mathit{p}}^{\mathit{sem}}+w_{\mathit{appe}}\cdot\mathit{s}_{\mathit{p}}^{\mathit{appe}}}{1+1}\cdot\mathit{conf}_{\mathit{p}}\text{,}$ (4)
where an appearance weight of $w_{\mathit{appe}}=1$ would compute the average but the baseline score is $w_{\mathit{appe}}=2$ based on some ablation experiments (see Appendix Sec. 6.1). The object matching scores of all the $N_{{\bm{I}}}^{\mathit{prop}}$ proposals, over all possible $N^{O}$ objects, are stored in the $N_{{\bm{I}}}^{\mathit{prop}}\times N^{O}$ instance score matrix. Note that, as small CT values squash down the appearance scores, a $w_{\mathit{appe}}=2$ is used.
@@
In the last step, we simply apply the Argmax function across the objects/rows of the instance score matrix. Each proposal gets assigned an object label and its object matching score as its corresponding confidence. Eventually, we obtain proposals consisting of: a bounding box of the object instance; its corresponding modal segmentation mask, which covers the visible instance part [hodan2024bopchallenge2023detection]; and an object label with a confidence score. To remove incorrectly labeled proposals and redundant ones, a confidence threshold filtering is applied with $\delta_{\mathit{conf}}=0.2$ followed by a Non-Maximum Suppression, respectively.